OpenStudy (anonymous):

Solve. 2x + 6 – 6x = 30

OpenStudy (anonymous):

A. –3 B. –6 C. 6 D. 3

OpenStudy (esshotwired):

x=-6

OpenStudy (anonymous):

are you sure

OpenStudy (esshotwired):

Positive

OpenStudy (esshotwired):

it would be B

OpenStudy (anonymous):

Solve. 4(x – 3) + 5x = –48 A. all real numbers B. 4 C. –4 D. 0

OpenStudy (anonymous):

2x + 6 – 6x = 30 (add 2x with -6x and subtract 6 from 30) -4x=24 x=-6

OpenStudy (anonymous):

oo ok thanks @Lynnetje

OpenStudy (anonymous):

what about the new one

OpenStudy (esshotwired):

x=-4

OpenStudy (anonymous):

wow a lot of people typing lol

OpenStudy (anonymous):

Solve. 6 – 3x = 4(3 – x) – 6 A. 0 B. all real numbers C. 6 D. –6

OpenStudy (anonymous):

2x+6-6x=30 add 2x to -6x and you get 4x and subtract 6 from 30 && you will get 24 -4x=24 divide by -4 && you should get x=-6

OpenStudy (esshotwired):

you would distribute so it would be 4x-12+5x=-48 then add 12 on both sides so it is 9x=-36 then divide by 9 and x=-4

OpenStudy (anonymous):

you need to start learning how to do this.

OpenStudy (anonymous):

cant just rely on people to answer your questions

OpenStudy (anonymous):

@JustTooMuchh what question are you on

OpenStudy (anonymous):

the first one .

OpenStudy (anonymous):

and @Lynnetje thats what this place is for

OpenStudy (anonymous):

can you help me wih this Solve. 6 – 3x = 4(3 – x) – 6 A. 0 B. all real numbers C. 6 D. –6

OpenStudy (esshotwired):

well i got up to 7x=24. My guess would probably be B

OpenStudy (anonymous):

ok how a bout this Frank’s quarter grades are 82.5, 94.7, and 87.9. What grade, rounded to the nearest percent, will Frank have to earn in the last quarter to receive at least a 89% final course grade? A. 90% B. 92% C. 91% D. 89%

OpenStudy (anonymous):

i am guessing a is that right

OpenStudy (anonymous):

or d

OpenStudy (esshotwired):

it would be C

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3 mor

OpenStudy (anonymous):

Which is the solution of the equation below, rounded to the nearest tenth? 2.1x – 4.09 = 6.98x + 1.73 A. –0.4 B. –0.5 C. –1.2 D. –1.1

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

i know i have been a user for ever i have another account ok

OpenStudy (esshotwired):

C, that is the last one I can help you with. I have my own work to do now.